OpenStudy (anonymous):

If Q(x)=X^2-5x+1, find (Q(1+h)-Q(1))/(h). Help me....

OpenStudy (anonymous):

difference quotient, lets take it one step at a time

OpenStudy (anonymous):

\[Q(x)=x^2-5x+1\] \[Q(x+h)=(x+h)^2-5(x+h)+1=x^2+2xh+h^2-5x-5x+1\]

OpenStudy (anonymous):

and then what's next? :| Thank You

OpenStudy (anonymous):

so \[Q(x+h)-Q(x)=x^2+2xh+x+h^2-5x-5h+1-(x^2-5x+1)\] \[=x^2+2xh+h^2-5x-5h+1-x^2+5x-1\] \[=\cancel{x^2}+2xh+h^2+\cancel{5x}+5h+\cancel{1}=2xh+h^2+5h\]

OpenStudy (anonymous):

you see that every term without an \(h\) in it is gone, now we can divide by \(h\)

OpenStudy (anonymous):

\[\frac{Q(x+h)-Q(x)}{h}=\frac{2xh+h^2+5h}{h}=\frac{h(2x+5+h)}{h}=2x+5+h\]

OpenStudy (anonymous):

oh damn damn damn it was \[\frac{Q(1+h)-Q(1)}{h}\] which is actually easier, replace all the \(x\)'s by 1 i can write it again if you like, or you can try it yourself, with \(1\) instead of \(x\) you will get \(2+5+h=7+h\)

OpenStudy (anonymous):

or impress you teacher, do it with \(x\) and then say "now we can replace \(x\) by any number we like to get the answer.